Determine how long 100$ must be left to accumulate at j12 = 18% for it to amount to twice the accumulated value of another $100 deposited at the same time at j2 = 10%. Ans; 8.55 years
Answer #1

100*(1+18%/12)^(12*t)=2*100*(1+10%/2)^(2*t)

=>12*t*log(1+18%/12)=log2+2*t*log(1+10%/2)

=>t*(12*log(1+18%/12)-2*log(1+10%/2))=log2

=>t=log(2)/(12*log(1+18%/12)-2*log(1+10%/2))

=>t=8.548 years